### v1

# An Integral Equation Arising from Time-Consistent Portfolio Selection

Abstract
This paper discusses a nonlinear integral equation arising from a class of time-consistent portfolio selection problem. We propose a unified framework requiring minimal assumptions, such as right-continuity of market coefficients and square-integrability of the market price of risk. Our main contribution is proving the existence and uniqueness of the square-integrable solution for the integral equation under mild conditions. Illustrative applications include the mean-variance portfolio selection and the utility maximization with random risk aversion.
Keywords: Integral Equation; Time-Inconsistency; Portfolio Selection; Equilibrium Strategy; Mean-Variance; Random Risk Aversion
1 Introduction
A time-inconsistent preference usually leads to the time-inconsistency of the optimal strategy: the strategy that is optimal today may not remain optimal tomorrow. To address the issue of time inconsistency, the intra-personal equilibrium, as a time-consistent solution to the problems with time-inconsistent preferences, is originally introduced by Strotz (1955). Since a precise definition of continuous-time intra-personal equilibrium was first provided by Ekeland and Lazrak (2006), the continuous-time intra-personal equilibrium has been studied for various time-inconsistent preferences. For instance, the mean-variance (MV) portfolio selection problem is investigated by Basak and Chabakauri (2010), Hu, Jin, and Zhou (2012, 2017), Björk, Murgoci, and Zhou (2014), and Dai, Jin, Kou, and Xu (2021); the non-exponential discounting problem is examined by Ekeland and Lazrak (2006), Ekeland and Pirvu (2008), Hamaguchi (2021), and Mbodji and Pirvu (2023). For a broader discussion on continuous-time intra-personal equilibrium, see Yan and Yong (2019), He and Zhou (2022), Hernández and Possamaï (2023), and the references therein.
The continuous-time intra-personal equilibria for a class of time-inconsistent preferences can be characterized by the solutions to a class of nonlinear integral (or differential) equations. Below, we provide three specific preferences to illustrate this. The first one is the MV preference with the risk aversion parameter being inversely proportional to the wealth level, which is proposed by Björk, Murgoci, and Zhou (2014). In a market with constant coefficients, the authors apply the Arzelà-Ascoli theorem to establish the existence and uniqueness of the solution to their integral equation (4.11), thereby confirm the existence and uniqueness of the equilibrium strategy among state-independent strategies. The second one is the preference with random risk aversion of Desmettre and Steffensen (2023), which addresses the problem of maximizing utility with random risk aversion. This preference is time-inconsistent due to the inclusion of nonlinear functions of expectation. The authors characterize the equilibrium strategies by an infinite-dimensional system of ODEs. Specifically, when the risk aversion has two possible values, this system reduces to a three-dimensional ODE system. However, the existence and uniqueness of the solution to this system is not provided there. The third one is the generalized disappointment aversion (GDA) preference, which is discussed in our recent work Liang, Wang, Xia, and Yuan (2024). Under GDA, the implicit definition of certainty equivalence introduces time inconsistency. We characterize the equilibrium strategies by a nonlinear integral equation.
This paper presents a unified approach to discuss the integral equation arising from a large class of time-consistent portfolio selection problems with portfolio bounds. Under mild conditions, we prove that this equation has a unique square-integrable solution.
Our main contribution is a unified framework that minimizes assumptions on market parameters. Specifically, we require only that the market coefficients are right-continuous and the market price of risk is square-integrable. In contrast, Björk, Murgoci, and Zhou (2014) and Desmettre and Steffensen (2023) assume constant market coefficients: the method in Björk, Murgoci, and Zhou (2014) cannot handle right-continuous coefficients, while Desmettre and Steffensen (2023) does not establish the existence and uniqueness of the solution to the ODE system. Liang, Wang, Xia, and Yuan (2024) assume that the market coefficients are right-continuous and bounded. Moreover, our equation can be applied to study the time monotonicity of equilibrium strategies. For instance, in the random risk aversion case discussed in Desmettre and Steffensen (2023), we theoretically demonstrate that in a single-stock market with constant market coefficients, the equilibrium investment decreases over time, exhibiting the so-called horizon effect.
The remainder of the paper is organized as follows. Section 2 investigates the existence and the uniqueness of the solution of a specific type of integral equation. Section 3 derives this integral equation from a class of time-consistent portfolio selection problem. Finally, Section 4 provides two illustrative applications.
2 An Integral equation
Let $T>0$ be a constant and $d\geq 1$ be an integer. Given a fixed function $h:[0,T)\times[0,\infty)\times\mathbb{R}\to\mathbb{R}$ and a fixed $\mathbb{R}^{d}$-valued function $\lambda\in L^{2}((0,T),\mathbb{R}^{d})$, we consider the following fully non-linear integral equation: [^1] For a matrix $A$, $A^{\top}$ denotes the transpose of $A$. For a vector $z\in\mathbb{R}^{d}$, $|z|\triangleq\sqrt{z^{\top}z}$. For any $0\leq t_{1}<t_{2}\leq T$, let $L^{2}((t_{1},t_{2}),\mathbb{R}^{d})$ denote the set of $\mathbb{R}^{d}$-valued Lebesgue measurable functions $\phi$ such that $\|\phi\|_{L^{2}((t_{1},t_{2}),\mathbb{R}^{d})}\triangleq\left(\int_{t_{1}}^{t_{2}}|\phi(s)|^{2}\mathrm{d}s\right)^{\frac{1}{2}}<\infty$.
$a(t)=\mathcal{P}_{t}\left(h\left(t,\sqrt{\int_{t}^{T}|a(s)|^{2}\mathrm{d}s},\int_{t}^{T}a^{\top}(s)\lambda(s)\mathrm{d}s\right)\lambda(t)\right),\quad t\in(0,T),$ (2.1)
where $a:(0,T)\to\mathbb{R}^{d}$ is a function to be determined by the equation and, for every $t\in[0,T)$, $\mathcal{P}_{t}$ is the projection from $\mathbb{R}^{d}$ to a convex and closed set $\mathbb{U}_{t}\subseteq\mathbb{R}^{d}$ with $\mathbf{0}\in\mathbb{U}_{t}$, i.e.,
$\mathcal{P}_{t}(k)\triangleq\arg\min_{\tilde{k}\in\mathbb{U}_{t}}|\tilde{k}-k|^{2},\quad k\in\mathbb{R}^{d}.$ (2.2)
Definition 2.1.
The function $h$ is called locally Lipchitz if, for any $M,N\in(0,\infty)$, there is a constant $\ell>0$ such that $|h(t,x_{1},y_{1})-h(t,x_{2},y_{2})|\leq\ell(|x_{1}-x_{2}|+|y_{1}-y_{2}|)$ for all $t\in[0,T)$ and $(x_{1},y_{1})$, $(x_{2},y_{2})\in[0,M]\times[-N,N]$. In this case, $\ell$ is called a Lipchitz constant of $h$ on $[0,T)\times[0,M]\times[-N,N]$.
Definition 2.2.
The function $h$ is called locally bounded if $h$ is bounded on $[0,T)\times[0,M]\times[-N,N]$ for any $M,N\in(0,\infty)$.
Definition 2.3.
The function $h$ is called locally I-bounded if
$\int_{0}^{T}\sup_{(x,y)\in[0,M]\times[-N,N]}|\mathcal{P}_{t}(h(t,x,y)\lambda(t))|^{2}|\mathrm{d}t<\infty$
for any $M,N\in(0,\infty)$. The function $h$ is called I-bounded if the above inequality holds when $[0,M]\times[-N,N]$ is replaced by $[0,\infty)\times\mathbb{R}$.
Remark 2.4.
It is evident that if $h$ is (locally) bounded, then $h$ is (locally) I-bounded as well. Furthermore, if $\mathbb{U}_{t}\subset\tilde{\mathbb{U}}$ for some bounded set $\tilde{\mathbb{U}}\subset\mathbb{R}^{d}$ and for all $t\in[0,T)$, then $h$ must be I-bounded.
For any $\phi\in L^{2}((t_{1},t_{2}),\mathbb{R}^{d})$, we will frequently use the following notations:
$v_{\phi}(t_{1},t_{2})\triangleq\|\phi\|^{2}_{L^{2}((t_{1},t_{2}),\mathbb{R}^{d})}=\int_{t_{1}}^{t_{2}}|\phi(s)|^{2}\mathrm{d}s,\quad y_{\phi}(t_{1},t_{2})\triangleq\int_{t_{1}}^{t_{2}}\phi^{\top}(s)\lambda(s)\mathrm{d}s.$
When $t_{2}=T$, we simply write $v_{\phi}(t_{1})$ and $y_{\phi}(t_{1})$ for $v_{\phi}(t_{1},T)$ and $y_{\phi}(t_{1},T)$, respectively.
We begin with the following theorem on the uniqueness of the solution.
Theorem 2.5.
If $h$ is locally Lipschitz, then, for every $\tau\in[0,T)$, the equation
$a(t)=\mathcal{P}_{t}\left(h\left(t,\sqrt{v_{a}(t)},y_{a}(t)\right)\lambda(t)\right),\quad t\in(\tau,T)$ (2.3)
has at most one solution $a\in L^{2}((\tau,T),\mathbb{R}^{d})$.
Proof.
Suppose that $a^{(i)}\in L^{2}((\tau,T),\mathbb{R}^{d})$, $i=1,2$, are two solutions of (2.3). Let
$M_{1}\triangleq\max\left\{\sqrt{v_{a^{(1)}}(\tau)},\sqrt{v_{a^{(2)}}(\tau)}\right\}$
and let $\ell_{1}$ be a Lipschitz constant of $h$ on
$[0,T)\times[0,M_{1}]\times\left[-M_{1}\sqrt{v_{\lambda}(0)},M_{1}\sqrt{v_{\lambda}(0)}\right].$
Using the fact: $|\mathcal{P}_{t}(h_{1})-\mathcal{P}_{t}(h_{2})|\leq|h_{1}-h_{2}|$ for any $t\in(0,T)$ and $h_{1},h_{2}\in\mathbb{R}^{d}$, we have, for every $t_{1}\in[\tau,T)$ and every $t\in(t_{1},T)$, we have
$\left|a^{(1)}(t)-a^{(2)}(t)\right|$ $\leq\ell_{1}|\lambda(t)|\left[\left|\sqrt{v_{a^{(1)}}(t)}\!-\!\sqrt{v_{a^{(2)}}(t)}\right|\!+\!\left|y_{a^{(1)}}(t)\!-\!y_{a^{(2)}}(t)\right|\right]$
$\leq\ell_{1}|\lambda(t)|\left[\sqrt{v_{a^{(1)}-a^{(2)}}(t)}+\sqrt{v_{\lambda}(t)}\cdot\sqrt{v_{a^{(1)}-a^{(2)}}(t)}\right]$
$\leq\ell_{1}|\lambda(t)|\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}.$
Therefore, we get
$\begin{split}\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{1})}\quad\forall\,t_{1}\in[\tau,T).\end{split}$ (2.4)
Let $s_{1}=\inf\left\{t_{1}\in[\tau,T):\ell_{1}\sqrt{v_{\lambda}(t_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\leq\frac{1}{2}\right\}$. Then $a^{(1)}=a^{(2)}$ a.e. on $(s_{1},T)$. If $s_{1}=\tau$, then $a^{(1)}=a^{(2)}$ a.e. on $(\tau,T)$. Otherwise, similar to (2.4), we have
$\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})}\leq\ell_{1}\sqrt{v_{\lambda}(t_{2},s_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(t_{2},s_{1})}\quad\forall\,t_{2}\in[\tau,s_{1}).$
Let $s_{2}=\inf\left\{t_{2}\in[\tau,s_{1}):\ell_{1}\sqrt{v_{\lambda}(t_{2},s_{1})}\left(1+\sqrt{v_{\lambda}(0)}\right)\leq\frac{1}{2}\right\}$. Then $a^{(1)}=a^{(2)}$ a.e. on $(s_{2},s_{1})$. If $s_{2}=\tau$, then $a^{(1)}=a^{(2)}$ a.e. on $(\tau,T)$. Otherwise, we repeat such a procedure. Up to finitely many times, we can obtain that $a^{(1)}=a^{(2)}$ a.e. on $(\tau,T)$. In this case, $v_{a^{(1)}}=v_{a^{(2)}}$ and $y_{a^{(1)}}=y_{a^{(2)}}$ everywhere on $(\tau,T)$, which implies that $a^{(1)}=a^{(2)}$ everywhere on $(\tau,T)$. ∎
As regards the local existence of the solution, we have the following lemma.
Lemma 2.6.
If $h$ is locally I-bounded and locally Lipschitz, then there exists some $\tau^{*}\in[0,T)$ such that Equation (2.3) has a solution $a\in L^{2}((\tau,T),\mathbb{R}^{d})$ for every $\tau\in[\tau^{*},T)$.
Proof.
For every $\tau\in[0,T)$, let
$\mathcal{S}_{\tau}=\left\{a\in L^{2}((\tau,T),\mathbb{R}^{d}):v_{a}(\tau)\leq 1\right\}.$
Obviously, $\mathcal{S}_{\tau}$ is a closed subset of $L^{2}((\tau,T),\mathbb{R}^{d})$. For every $a\in\mathcal{S}_{\tau}$, let
$(\mathcal{T}a)_{t}\triangleq\mathcal{P}_{t}\left(h\left(t,\sqrt{v_{a}(t)},y_{a}(t)\right)\lambda(t)\right),\quad t\in(\tau,T).$
First, we show that $\mathcal{T}$ is a map from $\mathcal{S}_{\tau}$ to $\mathcal{S}_{\tau}$ for every $\tau\in[0,T)$ that is sufficiently close to $T$. Indeed, for every $a\in\mathcal{S}_{\tau}$, we have
$\sqrt{v_{a}(t)}\leq 1\text{ and }\left|y_{a}(t)\right|\leq\sqrt{v_{\lambda}(0)}\quad\forall t\in[\tau,T).$
Then
$v_{\mathcal{T}a}(\tau)\leq\int_{\tau}^{T}\sup_{(x,y)\in[0,1]\times[-\sqrt{v_{\lambda}(0)},\sqrt{v_{\lambda}(0)}]}|\mathcal{P}_{t}(h(t,x,y)\lambda(t))|^{2}\mathrm{d}t.$
As $h$ is locally I-bounded, we can choose $\hat{\tau}\in[0,T)$ such that $v_{\mathcal{T}a}(\hat{\tau})\leq 1$, then $\mathcal{T}$ is a map from $\mathcal{S}_{\tau}$ to $\mathcal{S}_{\tau}$ for every $\tau\in[\hat{\tau},T)$.
Next, we show that there exists some $\tau^{*}\in[\hat{\tau},T)$ such that $\mathcal{T}$ is a contraction on $\mathcal{S}_{\tau}$ for every $\tau\in[\tau^{*},T)$. Indeed, let $\ell$ be a Lipschitz constant of $h$ on $[0,T)\times[0,1]\times[-\sqrt{v_{\lambda}(0)},\sqrt{v_{\lambda}(0)}]$. Using a similar way to (2.4), for any $\tau\in[\hat{\tau},T)$ and $a^{(i)}\in\mathcal{S}_{\tau}$, $i=1,2$, we have
$\sqrt{v_{\mathcal{T}a^{(1)}-\mathcal{T}a^{(2)}}(\tau)}\leq\ell\sqrt{v_{\lambda}(\tau)}\left(1+\sqrt{v_{\lambda}(0)}\right)\sqrt{v_{a^{(1)}-a^{(2)}}(\tau)}.$
Choose a $\tau^{*}\in[\hat{\tau},T)$ such that $\ell\sqrt{v_{\lambda}(\tau^{*})}\left(1+\sqrt{v_{\lambda}(0)}\right)\leq\frac{1}{2}$, then $\mathcal{T}$ is a contraction on $\mathcal{S}_{\tau}$ for every $\tau\in[\tau^{*},T)$.
Therefore, for every $\tau\in[\tau^{*},T)$, $\mathcal{T}$ has a unique fixed point in $\mathcal{S}_{\tau}$, as such (2.3) has a solution in $L^{2}((\tau,T),\mathbb{R}^{d})$. ∎
Assuming further that $h$ is I-bounded, we have the global existence of the solution, as the following theorem shows.
Theorem 2.7.
If $h$ is I-bounded and locally Lipschitz, then Equation (2.1) has a unique solution $a\in L^{2}((0,T),\mathbb{R}^{d})$.
Proof.
Theorem 2.5 implies the following fact: if, for each $i=1,2$, $a^{(i)}\in L^{2}((\tau_{i},T),\mathbb{R}^{d})$ is a solution of (2.3) for $\tau=\tau_{i}\in[0,T)$ with $0\leq\tau_{2}\leq\tau_{1}<T$, then $a^{(1)}=a^{(2)}$ on $(\tau_{1},T)$. Let $(\tau,T)\subseteq(-\infty,T)$ be the maximal interval for the existence of solution [^2] We have shown that $a^{(1)}$ can be extended to $a^{(2)}$ uniquely. Thus we can define the maximal interval for the existence of solution. In addition, we define $\lambda(t)=\lambda(0)$, $\mathcal{P}_{t}=\mathcal{P}_{0}$ and $h(t,\cdot,\cdot)=h(0,\cdot,\cdot)$ for $t<0$, thus $\tau$ may belong to $[-\infty,0)$. and $a$ be the solution on $(\tau,T)$. Then we must have $\lim\limits_{t\downarrow\tau}v_{a}(t)=\infty$ if $\tau>-\infty$. Suppose, for the sake of contradiction, that $\tau\in(-\infty,T)$ and $v_{a}(\tau)<\infty$. Consider the following equation for $\tilde{a}\in L^{2}((\tau-\varepsilon_{0},\tau),\mathbb{R}^{d})$:
$\tilde{a}(t)=\mathcal{P}_{t}\left(h\left(t,\sqrt{v_{\tilde{a}}(t,\tau)+v_{a}(\tau)},y_{\tilde{a}}(t,\tau)+y_{a}(\tau)\right)\lambda(t)\right),\quad t\in(\tau-\varepsilon_{0},\tau),$ (2.5)
where $\varepsilon_{0}>0$ is a constant to be determined. Using the same approach as in the proof of Lemma 2.6, there is a $\varepsilon_{0}>0$ such that (2.5) admits a solution in $L^{2}((\tau-\varepsilon_{0},\tau),\mathbb{R}^{d})$. Thus $a$ can be extended as the solution of (2.1) with $\tau$ replaced by $\tau-\varepsilon_{0}$. This contradicts the assumption that $(\tau,T)$ is the maximal interval for the existence of solution.
Assume that $h$ is I-bounded and $(\tau,T)$ is the maximal interval for the existence of solution. We only need to show that $\tau<0$. Indeed, suppose on the contrary $\tau\geq 0$. Then
$\infty=\lim\limits_{t\downarrow\tau}v_{a}(t)\leq v_{a}(\tau)\leq\int_{\tau}^{T}\sup_{(x,y)\in[0,\infty)\times\mathbb{R}}|\mathcal{P}_{t}(h(t,x,y)\lambda(t))|^{2}|\mathrm{d}t<\infty,$
which is impossible. ∎
Remark 2.8.
Let $h$ be locally Lipschitz. In the case where $\lambda\in L^{\infty}((0,T),\mathbb{R}^{d})$, $h$ is bounded, and no portfolio bounds are imposed, we have established the global existence and uniqueness of the solution in $L^{\infty}((0,T),\mathbb{R}^{d})$; see Liang, Wang, Xia, and Yuan (2024, Theorem 3.5). In contrast, Theorem 2.7 only requires $\lambda\in L^{2}((0,T),\mathbb{R}^{d})$ and $h$ is I-bounded to guarantee the existence and uniqueness of the solution in $L^{2}((0,T),\mathbb{R}^{d})$.
3 Time-consistent portfolio selection
We derive Equation (2.1) from a class of time-consistent portfolio selection problem in this section. Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},0\leq t\leq T\}$ is the augmented natural filtration generated by a standard $d$-dimensional Brownian motion $\{B(t),0\leq t\leq T\}$. In addition, suppose $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$ are given by
$dS_{i}(t)=S_{i}(t)[\mu_{i}(t)\mathrm{d}t+\sigma_{i}(t)\mathrm{d}B(t)],\quad t\in[0,T],$
where the market coefficients $\mu:[0,T]\rightarrow\mathbb{R}^{d}$ and $\sigma:[0,T]\rightarrow\mathbb{R}^{d\times d}$ are right-continuous and deterministic, and $\sigma_{i}$ denotes the $i$-th row of $\sigma$. We always assume that $\int_{0}^{T}|\mu(t)|\mathrm{d}t+\sum_{i=1}^{d}\int_{t}^{T}|\sigma_{i}(t)|^{2}\mathrm{d}t<\infty$ and $\sigma(t)$ is invertible for every $t\in[0,T]$. Let $\lambda(t)=(\sigma(t))^{-1}\mu(t)$ be the market price of risk. We also always assume that $\lambda\in L^{2}((0,T),\mathbb{R}^{d})$.
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables. A trading strategy is a process $\pi=\{\pi_{t},t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that $\int_{0}^{T}|\pi^{\top}_{t}\mu(t)|\mathrm{d}t+\int_{0}^{T}|\sigma^{\top}(t)\pi_{t}|^{2}\mathrm{d}t<\infty$ a.s., where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth into the stocks at time $t$. The self-financing wealth process $\{W^{\pi}_{t},0\leq t\leq T\}$ associated with a trading strategy $\pi$ satisfies the following stochastic differential equation (SDE):
[…]
The time-$t$ performance of a trading strategy $\pi$ is $J(t,\pi)$. We assume that $J(t,\cdot)$ can be (equivalently) represented as a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$. Here are some examples that will be further discussed in the next section. [^3] Other examples include the MV preference for the log returns of Dai, Jin, Kou, and Xu (2021), the constant-relative-risk-aversion (CRRA) betweenness preference discussed by liang2023dynamic, and the GDA preference discussed by Liang, Wang, Xia, and Yuan (2024).
Example 3.1.
The MV preference functional with state-dependent risk aversion parameter of Björk, Murgoci, and Zhou (2014) is given by
$\mathbb{E}_{t}\left[W^{\pi}_{T}\right]-\frac{\gamma}{2W^{\pi}_{t}}\textup{Var}_{t}\left[W^{\pi}_{T}\right],$ (3.2)
where $\gamma>0$ is a risk aversion index. The performance functional (3.2) is obviously equivalent to
$J(t,\pi)=\mathbb{E}_{t}\left[{W^{\pi}_{T}\over W^{\pi}_{t}}\right]-\frac{\gamma}{2}\textup{Var}_{t}\left[{W^{\pi}_{T}\over W^{\pi}_{t}}\right].$ (3.3)
Example 3.2.
The preference functional with random risk aversion of Desmettre and Steffensen (2023) is given by
$\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{\pi}_{T}\right)\right]\right)\mathrm{d}\Gamma(\gamma),$ (3.4)
where $\Gamma$ is a distribution function on $(0,\infty)$ and $u^{\gamma}$ are the CRRA utility functions given by
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},\quad\gamma>0,\gamma\neq 1,x>0,\\ \log x,\quad\gamma=1,x>0.\end{cases}$
The performance functional (3.4) is obviously equivalent to
$J(t,\pi)=\int_{(0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}\over W^{\pi}_{t}}\right)\right]\right)\mathrm{d}\Gamma(\gamma).$ (3.5)
Let $\mathbb{U}\subset\mathbb{R}^{d}$ be a convex and closed set with $\mathbf{0}\in\mathbb{U}$. A trading strategy $\pi$ is called admissible if $\pi_{t}\in\mathbb{U}$ a.s. and the time-$t$ functional $J(t,\pi)$ is well defined for all $t\in[0,T)$. Let $\Pi$ denote the set of all admissible strategies.
In general, as the optimal strategy at time $t$, which maximizes $J(t,\pi)$ over $\pi\in\Pi$, is not necessarily optimal at time $s>t$, we consider the so-called equilibrium strategies. Hereafter, we always consider a fixed $\bar{\pi}\in\Pi$, which is a candidate equilibrium strategy. For any $t\in[0,T)$, $\varepsilon\in(0,T-t)$ and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{U})$, let $\bar{\pi}^{t,\varepsilon,k}$ be given by
$\bar{\pi}^{t,\varepsilon,k}_{s}\triangleq\left\{\begin{aligned} &k,&s\in[t,t+\varepsilon),\\ &\bar{\pi}_{s},&s\notin[t,t+\varepsilon).\end{aligned}\right.$
The strategy $\bar{\pi}^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$. Following Hu, Jin, and Zhou (2012, 2017) and Yan and Yong (2019), we introduce the definition of equilibrium strategies as follows.
Definition 3.3.
$\bar{\pi}$ is called an equilibrium strategy if, for any $t\in[0,T)$, we have
$\lim_{\varepsilon\downarrow 0}\operatorname*{ess\,sup\,}_{\varepsilon_{0}\in(0,\varepsilon)}\frac{J(t,\bar{\pi}^{t,\varepsilon_{0},k})-J(t,\bar{\pi})}{\varepsilon_{0}}\leq 0\quad\text{a.s. for all }k\in L^{\infty}(\mathcal{F}_{t},\mathbb{U}).$ (3.6)
Because all the market coefficients are deterministic, it is natural to conjecture that an equilibrium strategy $\bar{\pi}$ is in the form of
$\bar{\pi}_{s}=(\sigma^{\top}(s))^{-1}a(s),\quad s\in[0,T),$ (3.7)
where $a\in L^{2}([0,T),\mathbb{R}^{d})$ is a deterministic and right-continuous $\mathbb{R}^{d}$-valued function such that $a(s)\in\sigma^{\top}(s)\mathbb{U}$ for all $s\in[0,T)$. Denote by $\mathcal{D}$ the trading strategies in the form of (3.7). For any given $t\in[0,T)$, it is easy to see that
$\frac{W^{\bar{\pi}}_{T}}{W^{\bar{\pi}}_{t}}\sim\mathrm{LogNormal}\left(y_{a}(t)-{\frac{1}{2}}v_{a}(t),v_{a}(t)\right).$ (3.8)
As $J(t,\cdot)$ is a functional of the conditional distribution of ${W^{\pi}_{T}\over W^{\pi}_{t}}$ given $\mathcal{F}_{t}$, we have [^4] Here, we implicitly assume that $\mathcal{D}\subset\Pi$.$J(t,\bar{\pi})=g(t,v_{a}(t),y_{a}(t))$ for some $g:[0,T)\times[0,\infty)\times\mathbb{R}\to\mathbb{R}$.
Theorem 3.4.
Assume that $\mathcal{D}\subset\Pi$, $g\in C^{0,1,1}([0,T)\times[0,\infty)\times\mathbb{R}))$, and $g_{v}(t,v,y)<0$ for all $(t,v,y)\in[0,T)\times[0,\infty)\times\mathbb{R}$. Then $\bar{\pi}=(\sigma^{\top})^{-1}a\in\mathcal{D}$ is an equilibrium strategy if and only if
$a(t)=\mathcal{P}_{t}\left(-\frac{g_{y}(t,v_{a}(t),y_{a}(t))}{2g_{v}(t,v_{a}(t),y_{a}(t))}\lambda(t)\right),\quad t\in[0,T),$ (3.9)
where $\mathcal{P}_{t}$ is the projection from $\mathbb{R}^{d}$ to $\mathbb{U}_{t}\triangleq\sigma^{\top}(t)\mathbb{U}$ defined in (2.2).
Proof.
As $\mathcal{D}\subset\Pi$, we know that $J(t,\bar{\pi}^{t,\varepsilon,k})$ is well defined for all $t\in[0,T)$, $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{U})$, and $\varepsilon\in(0,T-t)$. Let $\bar{\pi}^{t,\varepsilon,k}=(\sigma^{\top})^{-1}\tilde{a}$, then
$v_{\tilde{a}}(t)=v_{a}(t)+\int_{t}^{t+\varepsilon}(|\sigma^{\top}(s)k|^{2}-|\sigma^{\top}(s)\bar{\pi}_{s}|^{2})\mathrm{d}s,\quad y_{\tilde{a}}(t)=y_{a}(t)+\int_{t}^{t+\varepsilon}\mu^{\top}(s)(k-\bar{\pi}_{s})\mathrm{d}s.$
As $k$ is $\mathcal{F}_{t}$-measurable, we have $J(t,\bar{\pi}^{t,\varepsilon,k})=g(t,v_{\tilde{a}}(t),y_{\tilde{a}}(t))$. Then the condition (3.6) is equivalent to
$0\geq$ $\lim_{\varepsilon\to 0}\frac{g(t,v_{\tilde{a}}(t),y_{\tilde{a}}(t))\!-\!g(t,v_{a}(t),y_{a}(t))}{\varepsilon}$
$=$ $g_{v}(t,v_{a}(t),y_{a}(t))|(\sigma^{\top}(t)k|^{2}-|\sigma^{\top}(t)\bar{\pi}_{t}|^{2})+g_{y}(t,v_{a}(t),y_{a}(t))\mu^{\top}(t)(k-\bar{\pi}_{t})\quad\forall k\in\mathbb{R}^{d},$
which is further equivalent to
$g_{v}(t,v_{a}(t),y_{a}(t))|(\tilde{k}|^{2}-|a(t)|^{2})+g_{y}(t,v_{a}(t),y_{a}(t))\lambda^{\top}(t)(\tilde{k}-a(t))\leq 0\quad\forall\,\tilde{k}\in\sigma^{\top}(t)\mathbb{U}.$
That is,
$a(t)=\arg\min_{\tilde{k}\in\sigma^{\top}(t)\mathbb{U}}\left\{|\tilde{k}|^{2}+\frac{g_{y}(t,v_{a}(t),y_{a}(t))}{g_{v}(t,v_{a}(t),y_{a}(t))}\lambda^{\top}(t)\tilde{k}\right\}=\mathcal{P}_{t}\left(-\frac{g_{y}(t,v_{a}(t),y_{a}(t))}{2g_{v}(t,v_{a}(t),y_{a}(t))}\lambda(t)\right).$
Then by definition, $\bar{\pi}$ is an equilibrium strategy if and only if the equality in (3.9) holds on $[0,T)$. The equality for $t=0$ can be removed from the integral equation because $a_{0}$ contributes nothing to the integrals and we can straightforwardly put $a(0)=\mathcal{P}_{0}\left(-\frac{g_{y}(0,v_{a}(0),y_{a}(0))}{2g_{v}(0,v_{a}(0),y_{a}(0))}\lambda(0)\right)$ to get the equality at $t=0$ once $\{a(t),t\in(0,T)\}$ solves (3.9). ∎
Let
$h(t,x,y)=-\frac{g_{y}(t,x^{2},y)}{2g_{v}(t,x^{2},y)},\quad(t,x,y)\in[0,T)\times[0,\infty)\times\mathbb{R}.$ (3.10)
Then Equation (3.9) is exactly Equation (2.1). So we finish deriving Equation (2.1).
In the next section, we will apply the results on Equation (2.1) to some specific examples.
4 Applications
4.1 Dynamic Mean-Variance portfolio selection
In this subsection, we examine the problem that was studied by Björk, Murgoci, and Zhou (2014), [^5] The trading strategies in Björk, Murgoci, and Zhou (2014) is defined as the dollar amounts put into the risky assets. Their equilibrium strategy is formally a proportion strategy, but any perturbation of this strategy, which is a dollar-amount strategy, can lead to negative wealth. Therefore, as He and Jiang (2021, Page 3872) pointed out, the definition of the equilibrium strategy in Björk, Murgoci, and Zhou (2014) needs a modification such that their solution is indeed an equilibrium strategy. Our model, however, uses the proportion strategies to avoid this issue. where the MV portfolio selection was investigated under the assumption that the agent’s risk aversion is inversely proportional to the wealth level. Specifically, the agent’s MV preference functional at time $t$ can be equivalently given by (3.3). It is easy to see that $g$ and $h$ are independent of $t$ and, for all $(x,y)\in[0,\infty)\times\mathbb{R}$,
$g(v,y)=\mathrm{e}^{y}-\frac{\gamma}{2}\left(\mathrm{e}^{v+2y}-\mathrm{e}^{2y}\right),$ $g_{y}(v,y)=\mathrm{e}^{y}-\gamma\left(\mathrm{e}^{v+2y}-\mathrm{e}^{2y}\right),$
$g_{v}(v,y)=-\frac{\gamma}{2}\mathrm{e}^{v+2y},$ $h(x,y)=\frac{1+\gamma\mathrm{e}^{y}}{\gamma\mathrm{e}^{y+x^{2}}}-1.$ (4.1)
The above function $h$ is obviously unbounded: $\lim\limits_{y\to-\infty}h(x,y)=\infty$ for all $x\in[0,\infty)$. Therefore, we only have a local existence of the solution by Lemma 2.6. However, we can address this unboundedness issue by a prior estimation for the solutions in the next lemma.
Lemma 4.1.
Let function $h$ be given by (4.1). Suppose that $a\in L^{2}((0,T),\mathbb{R}^{d})$ is a solution to Equation (2.1). Then there exists a constant $M>0$, which depends only on $\lambda(\cdot)$, $T$, and $\gamma$, such that $h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\in[-M,1]$ for any $t\in[0,T]$.
Proof.
Obviously, $y_{a}$ satisfies the following equation
$y_{a}^{\prime}(t)=-\lambda^{\top}(t)\mathcal{P}_{t}\left(h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\lambda(t)\right),\quad t\in[0,T),\quad y_{a}(T)=0.$
As $\textbf{0}\in\mathbb{U}$, we have
$\left|\mathcal{P}_{t}\left(h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\lambda(t)\right)\right|\leq\left|h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\right|\cdot|\lambda(t)|.$
Then
$|y_{a}^{\prime}(t)|\leq\left|h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\right|\cdot|\lambda(t)|^{2}\leq\left(2+\frac{1}{\gamma}\mathrm{e}^{-v_{a}(t)-y_{a}(t)}\right)|\lambda(t)|^{2}.$ (4.2)
Next, observe that
$-v_{a}(t)-y_{a}(t)$ $\leq-v_{a}(t)+\sqrt{v_{\lambda}(0)}\sqrt{v_{a}(t)}$
$=\sqrt{v_{a}(t)}(\sqrt{v_{\lambda}(0)}-\sqrt{v_{a}(t)})\leq\frac{{v_{\lambda}(0)}}{4}.$
Combining this with (4.2) yields
$|y^{\prime}_{a}(t)|\leq\left(2+\frac{1}{\gamma}\mathrm{e}^{\frac{v_{\lambda}(0)}{4}}\right)|\lambda(t)|^{2}\triangleq C_{1}|\lambda(t)|^{2}.$
Then we obtain $|y_{a}(t)|\leq C_{1}v_{\lambda}(0)$ for all $t\in[0,T]$. Thus $h\left(\sqrt{v_{a}(t)},y_{a}(t)\right)\in[-M,1]$ for any $t\in[0,T]$ for some $M=M(T,\gamma,\lambda(\cdot))>0$. ∎
Using Lemma 4.1, we get the following proposition.
Proposition 4.2.
Let function $h$ be given by (4.1). Then Equation (2.1) has a unique solution in $L^{2}((0,T),\mathbb{R}^{d})$. Therefore, there is a unique equilibrium in $\mathcal{D}$.
Proof.
[…]
$\tilde{h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\in[-M,1],\quad t\in[0,T].$
Therefore $\tilde{h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))=h(v_{\tilde{a}}(t),y_{\tilde{a}}(t))$ and
$\tilde{a}(t)=\mathcal{P}_{t}\left(\tilde{h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right)=\mathcal{P}_{t}\left({h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right),\quad t\in[0,T).$
[…]
Remark 4.3.
In Björk, Murgoci, and Zhou (2014), an iterative approach and the Arzelà-Ascoli theorem are used to study an integral equation similar to (2.1) with constant market coefficients and without portfolio constraints. In contrast, we allow for time-dependent and discontinuous market coefficients and portfolio constraints, our equilibrium strategies are only right-continuous. Their method does not apply to our case of discontinuous market coefficients because their use of the Arzelà-Ascoli theorem requires that the market coefficients are continuously differentiable.
4.2 Equilibrium investment with random risk aversion
In this section, we consider the portfolio selection problem with random risk aversion that was studied in Desmettre and Steffensen (2023). The time-$t$ performance functional is given by (3.5). Let $\bm{R}$ be a random variable whose distribution function is $\Gamma$. Direct computation shows that $g$ is independent of $t$ and, for any $(v,y)\in[0,\infty)\times\mathbb{R}$, we have
$g(v,y)=g_{y}(v,y)=\int_{(0,\infty)}\mathrm{e}^{-\frac{1}{2}(\gamma v-2y)}\mathrm{d}\Gamma(\gamma)=\mathbb{E}\left[e^{-\frac{1}{2}(\bm{R}v-2y)}\right],$
$g_{v}(v,y)=-\frac{1}{2}\int_{(0,\infty)}\gamma\mathrm{e}^{-\frac{1}{2}(\gamma v-2y)}\mathrm{d}\Gamma(\gamma)=-\frac{1}{2}\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}(\bm{R}v-2y)}\right].$
Then, we know that $h$ is independent of $t$ and $y$, and has the following form:
$h(x)=\frac{\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]}{\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]},\quad x\in[0,\infty).$ (4.3)
To study the Lipschitz continuity of $h$, we introduce the following function:
$f(x)=\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right],\quad x\in[0,\infty).$
The next lemma summarizes some well-known properties of $f$ and $h$; see, Klenke (2020, Example 6.29).
Lemma 4.4.
We have the following conclusions:
  1. (i)
    $f,h\in C([0,\infty))\cap C^{\infty}((0,\infty))$.
  2. (ii)
    If $\mathbb{E}\left[\bm{R}\right]<\infty$, then $f\in C^{1}([0,\infty))$ and $f^{\prime}(x)=-\frac{1}{2}\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right].$
  3. (iii)
    If $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$, then $f\in C^{2}([0,\infty))$, $f^{\prime\prime}(x)=\frac{1}{4}\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]$, and $h\in C^{1}([0,\infty))$.
Combining Lemma 4.4 with Theorem 2.7 leads to the following proposition.
Proposition 4.5.
If $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and $h$ is bounded, then equilibrium exists uniquely in $\mathcal{D}$.
Remark 4.6.
In the case when $\Gamma$ is a two-point distribution, Desmettre and Steffensen (2023) characterize the equilibrium by a three-dimensional ODE system666 In fact, with appropriate transformations, their ODE system can be converted into our integral equation. without establishing the existence and uniqueness of the solution. Our approach, however, only requires that $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and that $h$ is bounded to establish the existence and uniqueness of the solution. The boundedness of $h$ is fulfilled when the support of $\Gamma$ has a positive lower bound, that is, $\Gamma(\gamma_{0})=0$ for some $\gamma_{0}>0$. Empirical studies, in fact, indicate that the risk aversion coefficient is bounded away from 00, as discussed in Chetty (2006) and Gandelman and Hernández-Murillo (2015), and the references therein.
Theorem 4.7.
Suppose that $d=1$, $\lambda(\cdot)>0$ is constant, and there are no portfolio bounds. Then the solution $a$ of (2.1) decreases with $t$. Furthermore, if $\Gamma$ is not a single-point distribution, $a$ is strictly decreasing with $t$.
Proof.
In this case, we know that
$a(t)=\frac{\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}{\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}\lambda.$
Then
$a’(t)=-\frac{1}{2}v^{\prime}(t)\frac{\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}-\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}{\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}}\lambda.$
Note that
$\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}$ $=\left(\mathbb{E}\left[\sqrt{\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}}\sqrt{\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}}\right]\right)^{2}$
$\leq\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right],$
[…]
$\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}=\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right],$
which is equivalent to $\Gamma$ being a single-point distribution. Therefore, $a$ is strictly decreasing with $t$ unless $\Gamma$ is a single-point distribution. ∎
Remark 4.8.
Suppose that $d=1$, $\lambda(\cdot)>0$ and $\sigma(\cdot)>0$ are constants, and there are no portfolio bounds. Then the equilibrium strategy $\bar{\pi}=\frac{1}{\sigma}a$ decreases with $t$. This indicates that the equilibrium strategy exhibits a horizon effect, which confirms the numerical results in Desmettre and Steffensen (2023).